johns garden is twice as long as it is wide. if the perimeter of the garden is 42 feet what is an equation to find the width of the garden
Answer by arKed(27)

P = 42
2W + 2L = P
where L = 2W
2W + 2 (2W) = P
2W +4W = P
6W / 6 = 42 / 6
W = 7.....ANS.
AL' SOL'N :
2W = P - 2L
2W = 42 - 2 (2W)
2W / 2= (42 - 2 (W))/ 2
W = 21 - 2W
3W /3 = 21 / 3
W = 7.....ANS.